Point Q is on line segment \overline{PR}
PR
. Given PQ=x,PQ=x, QR=6,QR=6, and PR=2x-10,PR=2x−10, determine the numerical length of \overline{PQ}.
PQ

.
1
Expert's answer
2020-09-29T18:59:40-0400

Here given that,

PQ=x,QR=6,PR=2x-10.

By the following diagram we can visualise that,

PR=PQ+QR



So,PR=PQ+QR

2X-10=x+6 (put the value of PQ,QR,PR from question)

2x-x=10+6

x=16

PQ=x=16 unit

From above the length of segment PQ is 16 unit.